Use the quadratic formula to solve
2x squared - 5x = 3

3x squared - 2x + 1 = 0

Answer by MathmaticsMaster(2)

2x^2 - 5x = 3 Original problem.
2x^2 - 5x - 3 = 0 Bring eveything to one side and remember that when you bring something over to another side, the sign changes.
(2x - 1)(x + 3) = 0 Factor completely.
2x - 1 = 0 x + 3 = 0 Finish solving.
2x = 1 x = -3 Finish solving.
x = 1/2 x = -3 <<<<< answers